OpenStudy (anonymous):

How to find if a line is above, below or on the plane, if line and plane are parallel to each other.

OpenStudy (shubhamsrg):

even if you dont know any formulae, just randomly pick any 2 points from the line ,find the distance of those 2 points from the place if distances are equal,then line is parallel to plane

OpenStudy (dumbcow):

lets try an example: plane \[2x-y-z = 3\] line \[x(t) = t +A\] \[y(t) = t+B\] \[z(t) = t+C\] if 2A -B-C = 3 , line is on plane 2A-B-C > 3 , line is below plane 2A-B-C < 3 , line is above plane

OpenStudy (anonymous):

^ There's no such thing as a line being above/below a plane, a line SEGMENT does, however.

OpenStudy (dumbcow):

a line can be above or below a plane as long as they are parallel

OpenStudy (anonymous):

Given a line (l): x=at+a0, y=bt+b0, z=ct+c0, and a plane (P): mx+ny+pz+q=0, the line can either (i) intersect the plane (ii) parallel to the plane (iii) lie on the plane. (i) happens iff am+bn+cp does not equal to 0. Otherwise, pick a point on the line, if it's also on the plane, then the line is on the plane, if not, then the line is parallel to the plane.

OpenStudy (dumbcow):

did that help? here is a more general explanation: plane \[Ax+By+Cz = D\] parallel line: \[x(t) = \frac{2}{A}t + e\] \[y(t) = -\frac{1}{B}t + f\] \[z(t) = -\frac{1}{C}t + g\] line is on plane if: \[Ae+Bf+Cg = D\] below plane if: \[Ae +Bf +Cg > D\] above plane if: \[Ae +Bf +Cg < D\]

OpenStudy (anonymous):

My bad,even with a line segment there's no such thing as below/above a plane. For example, given the plane x=1 and the line x=y=0, z=t.